# MoVieS: Motion-Aware 4D Dynamic View Synthesis in One Second

### v1

@@ 3.2.2 Prediction Heads
Shared aggregated video tokens from the feature backbone are fed into three parallel prediction heads that estimate the appearance, geometry, and motion of the dynamic scene, respectively. Each head adopts a DPT-style architecture (ranftl2021vision, ) to convert image tokens into dense predictions matching the input resolution, thereby producing dynamic splatter pixels.
@@
Different from previous feed-forward 3DGS reconstruction methods (charatan2024pixelsplat, ; chen2024mvsplat, ; zhang2024gs, ) that use a single head to predict all splatter pixel attributes, we adopt a decoupled design to better leverage geometric priors from the pretrained VGGT (wang2025vggt, ). Specifically, a dedicated depth head, initialized from VGGT, is used for geometry prediction to provide spatial grounding for splatter pixel construction, while a separate DPT as splatter head is trained from scratch for appearance rendering. We further incorporate a direct RGB shortcut (ye2024no, ) from the input image to the final convolution layer of the splatter head to preserve high-frequency details and enhance color fidelity.
@@
To capture scene dynamics, a novel motion head as shown in Figure 2 is introduced to predict dense deformation $\mathbf{m}(t)$ for each dynamic splatter pixel at any target moment. Temporal variation is enabled by injecting the sinusoidally encoded query time $t_{q}$ into the image tokens via adaptive layer normalization (xu2019understanding, ) before applying DPT convolutions. For each input frame $\mathbf{I}_{i}$ at time $t_{i}$, the motion head predicts its 3D movements toward $t_{q}$ in a shared world coordinate system. Thus, for $N$ input frames and $M$ query timestamps, this yields a motion map of shape $M\times N\times 3\times H\times W$.
@@ 3.3 Training
Table 1: Training Datasets. Nine datasets from diverse sources are utilized to train MoVieS at scale. “#Repeat” denotes dataset duplication count during integration to balance their contributions.
Dataset Dynamic? Depth? Tracking? Real? #Scenes #Frames #Repeat
RealEstate10K (zhou2018stereo, ) 70K 6.36M 1$\times$
TartanAir (wang2020tartanair, ) 0.4K 0.49M 100$\times$
MatrixCity (li2023matrixcity, ) 4.5K 0.31M 10$\times$
PointOdyssey (zheng2023pointodyssey, ) 0.1K 0.18M 1000$\times$
DynamicReplica karaev2023dynamicstereo 0.5K 0.26M 100$\times$
Spring (mehl2023spring, ) 0.03K 0.003M 2000$\times$
VKITTI2 (cabon2020virtual, ) 0.1K 0.03M 500$\times$
Stereo4D (jin2024stereo4d, ) 98K 19.6M 1$\times$
@@ 3.3.1 Dataset Curation
An ideal dataset for dynamic scene reconstruction would include synchronized multi-view videos with dense depth and point tracking annotations. However, such data is infeasible to capture and annotate at scale in practice. Instead, we leverage a diverse set of open-source datasets (zhou2018stereo, ; wang2020tartanair, ; li2023matrixcity, ; mehl2023spring, ; cabon2020virtual, ; zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ), each providing complementary supervision, as shown in Table 1. With the flexible model design, MoVieS can be jointly trained on these heterogeneous sources by aligning objectives to their respective annotations. More details about data curation are provided in Appendix A.
@@ 3.3.2 Objectives
Depth loss is computed as the mean squared error (MSE) between the predicted and ground truth depth maps, along with their spatial gradients, after filtering out invalid values. Rendering loss combines pixel-wise MSE and perceptual loss (zhang2018perceptual, ) between 3DGS-rendered images under corresponding camera views and the video frames at target timestamps.
@@
$\ =\frac{1}{P}\sum_{i\in\Omega}\lambda_{\text{pt}}\|\Delta\hat{\mathbf{x}}_{i}-\Delta\mathbf{x}_{i}\|_{1}+\frac{1}{P^{2}}\sum_{(i,j)\in\Omega\times\Omega}\lambda_{\text{dist}}\|\Delta\hat{\mathbf{x}}_{i}\cdot\Delta\hat{\mathbf{x}}_{j}^{\top}-\Delta\mathbf{x}_{i}\cdot\Delta\mathbf{x}_{j}^{\top}\|_{1},$ (4)
where $\Omega$ denotes the set of all valid $P$ tracked points, and $\Omega\times\Omega$ is its Cartesian product. Ablation study on the effectiveness of these two types of motion supervision is presented in Section 4.4.
@@
Similar to VGGT (wang2025vggt, ), we normalize the 3D scene scale by the average Euclidean distance from each 3D point to the origin of the canonical world coordinate system. As a result, unlike some other reconstruction methods (wang2024dust3r, ; leroy2024grounding, ; wang2025continuous, ), we do not apply additional normalization in the depth or motion loss. We also omit confidence-aware weighting (wang2024dust3r, ; wang2025vggt, ) for simplicity.
@@ 4 Experiments
4.1 Experimental Settings
Implementation
MoVieS is built on a geometrically pretrained transformer, VGGT (wang2025vggt, ), with splatter head and camera/time embeddings trained from scratch. AdamW (loshchilov2017decoupled, ) with cosine learning rate scheduling and linear warm-up is used for optimization. We observed that training MoVieS is particularly unstable, likely arising from sparse annotations and the heterogeneous nature of the training data. A curriculum strategy is employed that gradually increases the training complexity, involving (1) pretraining on static scenes, (2) dynamic scenes with varying views, and (3) fine-tuning on high resolution. Several techniques, such as gsplat (ye2025gsplat, ) rendering backend, DeepSpeed (rasley2020deepspeed, ), gradient checkpointing (chen2016training, ), gradient accumulation, and bf16 mixed precision, are adopted to improve memory and computation efficiency. Training completes in approximately 5 days using 32 H20 GPUs. Please refer to Appendix A for more details about implementations.
Evaluation
We evaluate the capabilities of MoVieS on two primary tasks: novel view synthesis (Section 4.2) and 3D point tracking (Section 4.3). Following prior works (charatan2024pixelsplat, ; chen2024mvsplat, ; xu2024depthsplat, ; wang2024shape, ; sun2024splatter, ; lei2025mosca, ; liang2024feed, ), RealEstate10K (zhou2018stereo, ) is used to evaluate novel view synthesis performance on static scenes. DyCheck (gao2022dynamic, ) and NVIDIA dynamic scene dataset (yoon2020novel, ) are adopted for dynamic scenes, in terms of PSNR, SSIM and LPIPS (zhang2018perceptual, ). Since the DyCheck dataset contains invisible regions in the target novel views, we compute reconstruction metrics (denoted by a prefix “m”) using the provided covisibility masks, ensuring a fair comparison across methods. For 3D point tracking, the TAPVid-3D (koppula2024tapvid, ) benchmark serves as the evaluation protocol, covering both indoor and outdoor scenes across three datasets. End-point Euclidean error in the 3D space (EPE3D) and the percentage of 3D points within $0.05$ and $0.10$ units of the ground-truth locations ($\delta^{0.05}_{3D}$ and $\delta^{0.10}_{3D}$) are utilized as metrics.
4.2 Novel View Synthesis
4.2.1 Static Scene View Synthesis
As a special case of dynamic scenes, we first evaluate MoVieS on a static dataset RealEstate10K (zhou2018stereo, ), comparing against several state-of-the-art feed-forward static scene reconstruction methods, including DepthSplat (xu2024depthsplat, ), our reimplementation of GS-LRM (zhang2024gs, ), and our method pretrained solely on the first-stage static reconstruction. As shown in Table 2, although MoVieS is primarily designed for dynamic scenes, it maintains competitive performance on static scenes. Notably, when processing static inputs, our predicted motion naturally converges to zero, demonstrating the ability of MoVieS to implicitly differentiate between static and dynamic regions without explicit supervision.
4.2.2 Dynamic Scene View Synthesis
Dynamic scene view synthesis is evaluated on two benchmarks: (1) DyCheck (gao2022dynamic, ), with 7 daily videos captured by 3 synchronized cameras, and (2) NVIDIA (yoon2020novel, ), containing 9 dynamic scenes recorded by a 12-camera static rig. For DyCheck, we sample 13 frames of size $518\times 518$ from a random clip from iPhone-record videos, and include the other two cameras for evaluation. For NVIDIA, 12 input views at $379\times 672$ resolution are sampled in a round-robin style (gao2021dynamic, ; liu2023robust, ; liang2024feed, ), where the i-th frame comes from the i-th camera at timestep i. We compare against three state-of-the-art optimization-based methods (wang2024shape, ; sun2024splatter, ; lei2025mosca, ) for dynamic 3D Gaussian reconstruction, and also report static feed-forward baselines for reference.
As shown in Table 2, MoVieS achieves competitive or superior performance compared to baselines, while requiring only 0.93s per scene, which is orders of magnitude faster than prior approaches that rely on heavy pretrained models and complex multi-stage pipelines. Qualitative visualizations in Figure 3 further highlight the strength of our approach. While MoSca (lei2025mosca, ) demonstrates impressive performance, it struggles with sparse inputs and often overfits to seen poses, producing spiky and over-saturated artifacts under novel views and timesteps. In contrast, MoVieS leverages large-scale learned priors to generalize more effectively, yielding smoother and more realistic results.
To ensure fair comparison and better reflect real-world scenarios, no video masks for dynamic objects are used in our experiments. It poses a major challenge for optimization-based methods like Shape-of-Motion (wang2024shape, ) and Splatter-a-Video (sun2024splatter, ), which rely heavily on explicit motion segmentation. The challenge is particularly evident on the NVIDIA dataset, where significant camera shake hinders disentangling scene dynamics, leading to notably degraded performance and even worse than static baselines. In contrast, our method exhibits strong robustness by directly learning to model motion.
Figure 3: Novel View Synthesis for Dynamic Scenes. Given a monocular video, we compare synthesized views from a novel view across different methods. Regions invisible in the input are rendered as black or white, depending on the rendering implementation. More results in Figure 6.
@@ 4.2.2 Dynamic Scene View Synthesis
Table 2: Evaluation on Novel View Synthesis. The best, second best and third best results are highlighted for clarity. $\dagger$ indicates our reimplemented version of GS-LRM (zhang2024gs, ).
@@
DepthSplat (xu2024depthsplat, ) 0.60s 26.57 80.06 0.124 13.83 0.4364 0.3850 17.16 0.5002 0.3023
GS-LRM (zhang2024gs, ) 0.57s 26.94 79.13 0.139 14.60 0.4535 0.3775 17.83 0.4988 0.3265
Ours (static) 0.84s 27.60 81.25 0.113 15.24 0.4784 0.3783 18.73 0.5042 0.2959
@@
Splatter-a-Video (sun2024splatter, ) 37min - - - 13.61 0.3131 0.5706 14.39 0.2538 0.5983
Shape-of-Motion (wang2024shape, ) 10min - - - 17.96 0.5662 0.3463 15.30 0.3169 0.5087
MoSca (lei2025mosca, ) 45min - - - 18.24 0.5514 0.3698 21.45 0.7123 0.2653
Ours 0.93s 26.98 81.75 0.111 18.46 0.5887 0.3094 19.16 0.5141 0.3152
4.3 3D Point Tracking
Trained on large-scale point tracking datasets (zheng2023pointodyssey, ; karaev2023dynamicstereo, ; jin2024stereo4d, ), the proposed method can also densely track any 3D point corresponding to a pixel across video frames (see Figure 7 in the appendix). We compare MoVieS against three strong baselines, including two state-of-the-art 2D point tracking methods, BootsTAP (doersch2024bootstap, ) and CoTracker3 (karaev2024cotracker, ), as well as a native 3D point tracking approach, SpatialTracker (xiao2024spatialtracker, ). For 2D trackers, a recent video depth estimation model (chen2025video, ) and ground-truth camera intrinsics are used to unproject tracked points into 3D space. To account for scale differences across methods, we normalize all predicted 3D points by their median norm before evaluation.
Quantitative results are reported in Table 3. While 3D-based SpatialTracker generally outperforms 2D-based approaches, all of them rely heavily on pretrained monocular depth estimators for geometry reasoning, introducing significant noise and inconsistency in the 3D space. In contrast, MoVieS directly estimates 3D point positions in a shared world coordinate, enabling more accurate and robust 3D tracking, and achieves consistently superior or competitive performance across all datasets.
Table 3: Evaluation on 3D Point Tracking. The best, second best and third best results are highlighted for clarity. $\dagger$ denotes 3D points obtained by combining a depth estimation model (chen2025video, ).
@@
$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$downarrow EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$downarrow EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
BootsTAPIR (doersch2024bootstap, ) 0.5539 17.73% 32.97% 0.0617 55.82% 75.66% 0.0650 69.28% 87.95%
CoTracker3 (karaev2024cotracker, ) 0.5614 19.88% 35.82% 0.0637 55.30% 77.55% 0.0617 69.27% 88.04%
SpatialTracker (xiao2024spatialtracker, ) 0.5413 18.08% 38.23% 0.0648 56.58% 80.67% 0.0519 72.91% 89.86%
@@ 4.4 Ablation and Analysis
$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
@@
$\uparrow$ mPSNR $\uparrow$ mSSIM $\downarrow$ mLPIPS $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$uparrow $\delta^{0.10}_{3D}$
NVS w/o motion 15.82 0.4796 0.3741 18.38 0.4856 0.3010 0.7938 19.58% 32.86%
Motion w/o NVS 16.26 0.4556 0.3461 18.98 0.4939 0.3207 0.3801 24.72% 42.92%
Ours 18.46 0.5887 0.3094 19.16 0.5041 0.3152 0.2153 52.05% 71.63%

### v2

@@ 3.2.2 Prediction Heads
Tokens from the feature backbone are fed into three parallel prediction heads that estimate the appearance, geometry and motion of a dynamic scene respectively, as shown in Figure 1. Each head adopts a DPT-style architecture [53] to convert aggregated tokens into dense predictions matching the input resolution and product dynamic splatter pixels.
@@
Different from previous feed-forward 3DGS reconstruction methods [5, 9, 92] that use a single head to predict all splatter pixel attributes, we adopt a decoupled design to better leverage geometric priors from the pretrained VGGT [68]. Specifically, a dedicated depth head, initialized from VGGT, is used for geometry prediction to provide spatial grounding for splatter pixel construction, while another DPT as splatter head is trained from scratch for appearance rendering. We further incorporate a direct RGB shortcut [88] from the input image to the final convolution layer of the splatter head to preserve high-frequency details and enhance color fidelity.
@@
To capture scene dynamics, a novel motion head as shown in Figure 2 is introduced to predict dense deformation $\mathbf{m}(t)$ for each dynamic splatter pixel at any target moment. Temporal variation is enabled by injecting the sinusoidally encoded query time $t_{q}$ into the aggregated tokens via adaptive layer normalization [81] before applying DPT convolutions. For each input frame $\mathbf{I}_{i}$ at time $t_{i}$, the motion head predicts its 3D movements $\Delta\mathbf{x}$ and Gaussian attribute deformation $\Delta\mathbf{a}$ toward $t_{q}$ in a shared world coordinate system. To visualize motion maps, the XYZ coordinates in 3D space are jointly normalized by their shared minimum and maximum values to the range $[0,1]$, and then mapped to RGB channels for visualization.
Table 1: Training Datasets. Eight datasets from diverse sources are utilized to train MoVieS at scale.
Dataset Dynamic? Depth? Tracking? Real? #Scenes
RealEstate10K [97] 70K
TartanAir [75] 0.4K
MatrixCity [29] 4.5K
PointOdyssey [96] 0.1K
DynamicReplica [21] 0.5K
Spring [44] 0.03K
VKITTI2 [3] 0.1K
Stereo4D [20] 98K
Figure 3: Novel View Synthesis for Dynamic Scenes. Given a monocular video, we compare synthesized novel views of different methods. Invisible regions are rendered as black or white, depending on the implementation. More results are in the supplementary material.
@@ 3.3.1 Dataset Curation
An ideal dataset for dynamic reconstruction would include synchronized multi-view videos with dense depth and point tracking annotations. However, such data is infeasible to capture and annotate at scale in practice. Instead, we leverage a diverse set of open-source datasets [97, 75, 29, 44, 3, 96, 21, 20], each providing complementary supervision, as shown in Table 1. With the flexible model design, MoVieS can be trained on these heterogeneous sources by aligning objectives to their respective annotations. For datasets without an official test split, we randomly sample 10% of the data as the test set for generalization evaluation. Details about data curation are in the supplementary material.
@@ 3.3.2 Objectives
Depth loss is computed as the mean squared error (MSE) between the predicted depth maps $\hat{D}_{i}$ and ground truth depth maps $D_{i}$, along with their spatial gradients, after filtering out invalid values.
$\mathcal{L}_{\text{depth}}\coloneq\sum_{i=1}^{N}\|D_{i}-\hat{D}_{i}\|_{2}+\|\nabla D_{i}-\nabla\hat{D}_{i}\|_{1}.$ (3)
Rendering loss combines pixel-wise MSE and perceptual loss [93] between 3DGS-rendered images $\hat{I}$ under $M$ camera views and the video frames at corresponding target timestamps, which are randomly sampled during training.
$\mathcal{L}_{\text{rendering}}\coloneq\sum_{v=1}^{M}\|I_{v}-\hat{I}_{v}\|_{2}+\lambda_{\text{LPIPS}}\cdot\text{LPIPS}(I_{v},\hat{I}_{v}).$ (4)
Weighting terms $\lambda_{\text{d}}$, $\lambda_{\text{r}}$ and $\lambda_{\text{LPIPS}}$ are set to $1$, 11 and $0.5$ by default respectively.
@@
$\ =\frac{1}{P}\sum_{i\in\Omega}\lambda_{\text{pt}}\|\Delta\hat{\mathbf{x}}_{i}-\Delta\mathbf{x}_{i}\|_{1}$ (6)
$\ +\frac{1}{P^{2}}\sum_{(i,j)\in\Omega\times\Omega}\lambda_{\text{dist}}\|\Delta\hat{\mathbf{x}}_{i}\cdot\Delta\hat{\mathbf{x}}_{j}^{\top}-\Delta\mathbf{x}_{i}\cdot\Delta\mathbf{x}_{j}^{\top}\|_{1},$ (7)
where $\Omega$ denotes the set of valid $P$ tracked points, and $\Omega\times\Omega$ is its Cartesian product. Weighting term $\lambda_{\text{m}}$, $\lambda_{\text{pt}}$, and $\lambda_{\text{dist}}$ are set to $10$, 11, and $10$ respectively. Ablation study on the effectiveness of these two types of motion supervision is presented in Sec. 4.4.
@@
Similar to VGGT [68], we normalize the 3D scene scale by the average Euclidean distance from each 3D point to the origin of the canonical world coordinate system. As a result, unlike some other reconstruction methods [74, 27, 73], we do not apply additional normalization in the depth or motion loss. We also omit confidence-aware weighting [74, 68] for simplicity and more stable training.
@@ 4 Experiments
Table 2: Evaluation on Novel View Synthesis. The best, second best and third best results are highlighted for clarity. $\dagger$ indicates our reimplemented version of GS-LRM [92]. “Ours (static)” refers to our method pretrained solely on static datasets without the motion head. The same camera parameters are provided for all methods for fair comparison.
@@
DepthSplat [79] 0.60s 26.57 80.06 0.124 13.83 43.64 0.3850 17.16 50.02 0.3023
GS-LRM [92] 0.57s 26.94 79.13 0.139 14.60 45.35 0.3775 17.83 49.88 0.3265
Ours (static) 0.84s 27.60 81.25 0.113 15.24 47.84 0.3783 18.73 50.42 0.2959
@@
Splatter-a-Video [62] 37min - - - 13.61 31.31 0.5706 14.39 25.38 0.5983
Shape-of-Motion [72] 10min - - - 17.96 56.62 0.3463 15.30 31.69 0.5087
MoSca [26] 45min - - - 18.24 55.14 0.3698 21.45 71.23 0.2653
Ours 0.93s 26.98 81.75 0.111 18.46 58.87 0.3094 19.16 51.41 0.3152
Table 3: Evaluation on 3D Point Tracking. The best, second best and third best results are highlighted for clarity. $\dagger$ denotes combining a depth estimation model [6]. All methods are given the same camera information for unprojecting 3D tracked points.
@@
$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$ $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$ $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$
BootsTAPIR [10] 0.5539 17.73% 32.97% 0.0617 55.82% 75.66% 0.0650 69.28% 87.95%
CoTracker3 [22] 0.5614 19.88% 35.82% 0.0637 55.30% 77.55% 0.0617 69.27% 88.04%
SpatialTracker [77] 0.5413 18.08% 38.23% 0.0648 56.58% 80.67% 0.0519 72.91% 89.86%
@@
$\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$
@@
$\uparrow$ mPSNR $\uparrow$ mSSIM $\downarrow$ mLPIPS $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ LPIPS $\downarrow$ EPE3D $\uparrow$ $\delta^{0.05}_{3D}$ $\uparrow$ $\delta^{0.10}_{3D}$
NVS w/o motion 15.82 47.96 0.3741 18.38 48.56 0.3010 0.7938 19.58% 32.86%
Motion w/o NVS 16.26 45.56 0.3461 18.98 49.39 0.3207 0.3801 24.72% 42.92%
Ours 18.46 58.87 0.3094 19.16 50.41 0.3152 0.2153 52.05% 71.63%
Figure 5: Zero-shot Applications. Predicted motion maps from our model can be directly applied to downstream tasks, such as (a) scene flow estimation and (b) moving object segmentation, in a zero-shot manner, without any task-specific fine-tuning or supervision.
4.1 Experimental Settings
Implementation
MoVieS is built on a geometrically pretrained transformer, VGGT [68], via full fine-tuning. The splatter head and camera/time embeddings are trained from scratch. These new embeddings are zero-initialized to provide a stable initialization without interfering with the pretrained model. AdamW [41] with cosine learning rate scheduling and linear warm-up is used for optimization. We observed that training MoVieS is particularly unstable, likely due to sparse annotations and the heterogeneous nature of training datasets. A curriculum strategy is used that gradually increases the training complexity, involving (1) pretraining on static scenes, (2) dynamic scenes with varying views, and (3) fine-tuning in high resolution. Several techniques, such as gsplat rendering backend [89], DeepSpeed [55], gradient checkpointing [7], gradient accumulation, and bf16 mixed precision, are adopted to improve memory and computation efficiency. Training completes in approximately 5 days using 32 H20 GPUs. More details are provided in the supplementary material.
Evaluation
We evaluate MoVieS on two primary tasks: novel view synthesis (Sec. 4.2) and 3D point tracking (Sec. 4.3). Considering pose-aware scenarios, we provide the same camera parameters for all methods to ensure a fair comparison. Following prior works [5, 9, 79, 72, 62, 26, 33], RealEstate10K [97] is used to evaluate novel view synthesis performance on static scenes. DyCheck [14] and NVIDIA dynamic scene dataset [90] are adopted for dynamic scenes, in terms of PSNR, SSIM and LPIPS [93]. Since the DyCheck dataset contains invisible regions in the target novel views, we compute reconstruction metrics using the provided covisibility masks (denoted by a prefix “m”), ensuring a fair comparison across methods. For 3D point tracking, the TAPVid-3D [25] benchmark serves as the evaluation protocol, covering both indoor and outdoor scenes across three datasets. End-point Euclidean error in the 3D space (EPE3D) and the percentage of 3D points within $0.05$ and $0.10$ units of the ground-truth locations ($\delta^{0.05}_{3D}$ and $\delta^{0.10}_{3D}$) are utilized as metrics. A large pretrained video depth estimation model, Video Depth Anything [6], and the same camera information are combined to unproject the predicted 2D tracking points of baselines into the 3D space for fair comparison.
4.2 Novel View Synthesis
4.2.1 Static Scene View Synthesis
As a special case of dynamic scenes, we first evaluate MoVieS on a static dataset, RealEstate10K [97], comparing against several state-of-the-art feed-forward static scene reconstruction methods, including DepthSplat [79], our reimplementation of GS-LRM [92], and our method pretrained solely on the first-stage static reconstruction. As shown in Table 2, although MoVieS is primarily designed for dynamic scenes, it maintains competitive performance on static scenes. Notably, when processing static inputs, our predicted motion naturally converges to zero (less than 1e-3), demonstrating the ability of MoVieS to implicitly differentiate between static and dynamic regions.
4.2.2 Dynamic Scene View Synthesis
We compare against three state-of-the-art open-sourced methods [72, 62, 26] for dynamic 3D Gaussian reconstruction, and also report static feed-forward baselines for reference. For DyCheck, we sample 13 frames of size $518\times 518$ from a random clip of 65 frames from iPhone-record videos, and include the other two cameras for evaluation. For NVIDIA, 12 input views at $379\times 672$ resolution are sampled in a round-robin style [13, 38, 33], where the i-th frame comes from the i-th camera at timestep i.
As shown in Table 2, MoVieS achieves competitive or superior performance compared to baselines, while requiring only 0.93s per scene, which is orders of magnitude faster than prior approaches that rely on heavy pretrained models and complex multi-stage pipelines. Qualitative visualizations in Figure 3 further highlight the strength of our approach. While MoSca [26] demonstrates impressive performance, it struggles with sparse inputs and often overfits to seen poses, producing spiky and over-saturated artifacts under novel views and timesteps. In contrast, MoVieS leverages large-scale learned priors to generalize more effectively, yielding smoother and more realistic results.
To ensure fair comparison and better reflect real scenarios, no video masks for dynamic objects are used in our experiments. It poses a major challenge for optimization-based methods like Shape-of-Motion [72] and Splatter-a-Video [62], which rely heavily on explicit motion segmentation. The challenge is particularly evident on the NVIDIA dataset, where significant camera shake hinders disentangling scene dynamics, leading to notably degraded performance and even worse than static baselines. In contrast, MoVieS exhibits strong robustness by directly learning to model motion. To evaluate the robustness of our method, visualization results on the in-the-wild DAVIS dataset [49] are provided in the supplementary material, where camera poses are estimated by MegaSaM [31].
4.3 3D Point Tracking
Trained on large-scale point tracking datasets [96, 21, 20], the proposed method can also densely track any 3D point corresponding to a pixel across video frames. We compare MoVieS against three strong baselines, including two state-of-the-art 2D point tracking methods, BootsTAP [10] and CoTracker3 [22], as well as a native 3D point tracking approach, SpatialTracker [77]. For 2D trackers, a recent video depth estimation model [6] and ground-truth camera intrinsics are used to unproject tracked points into 3D space. Considering scale differences across methods, we normalize all predicted 3D points by their median norm before evaluation.
Quantitative results are reported in Table 3. While 3D-based SpatialTracker generally outperforms 2D-based approaches, all of them rely heavily on pretrained monocular depth estimators for geometry reasoning, introducing significant noise and inconsistency in the 3D space. In contrast, MoVieS directly estimates 3D point positions in a shared world coordinate, enabling more accurate and robust 3D tracking, and achieves consistently superior or competitive performance across all datasets. Visualization results are provided in the supplementary material.
4.4 Ablation and Analysis